OpenStudy (anonymous):

Pilar spends a total of 60 hours per week at school and her job. She attends school from 8:45 AM until 3:45 PM, Monday through Friday. Which equation can be used to find t, the maximum number of hours Pilar works at her job each week?

OpenStudy (anonymous):

Please explain

OpenStudy (anonymous):

she spend 7 hours for five days in school. which is 35 hours and the remaining from 60 hours is for her work which is 25 hours. the equation is. 60- (7)(5) = t

OpenStudy (anonymous):

the answers are A.t=60-(5*7) B. t=60-(5*8) C.t=60-7+7+7+7+7 D.t=5*7-60

OpenStudy (anonymous):

okk, tnx.

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

im going to fan you, kumar.

OpenStudy (anonymous):

Its k, tnx for taking the time to help me.